# Aragog: Just-in-Time Model Routing for Scalable Serving of Agentic Workflows

### v1

@@ Abstract
Agentic workflows have emerged as a powerful paradigm for solving complex, multi-stage tasks, but serving them at scale is computationally expensive given the many LLM inferences that each request must pass through. Configuration selection, or the cost-aware assignment of workflow agents to specific LLMs, can reduce these costs, but existing approaches bind configuration decisions before request execution, making them ill-suited for the heterogeneous and lengthy execution of workflows. Specifically, system loads can fluctuate rapidly and substantially during a request’s lifetime, causing fixed configurations to quickly become suboptimal. We present Aragog, a system that progressively adapts a request’s configuration throughout its execution to match runtime dynamics. To make this practical despite the massive space of workflow configurations, Aragog decouples the problem into two core elements – a one-time routing step that identifies all accuracy-preserving configurations, and a cheap per-stage scheduler that selects among them using up-to-date system observations – and introduces novel strategies to accelerate each. Across diverse workflows and model families, Aragog increases maximum serving throughput by 50.0–217.0% and reduces median latency by 32.5–78.9% at peak request rates, while maintaining accuracy comparable to the most expensive configurations.
@@ 1 Introduction
Autonomous agents, or Large Language Models (LLMs) augmented with tools, memory, and planning, have shown impressive ability to tackle complex tasks autonomously [yang2024sweagent, snowflake-agentic-workflow]. However, fully agentic systems often lack the structure and reliability needed in production [multiagentllmsystemsfail]. This has led to the rise of agentic workflows: directed graphs that orchestrate multiple agents [langchain, dspy, autogen, metagpt], each responsible for a well-defined subtask and passing outputs downstream. These workflows preserve the generative strengths of agents while offering the observability and debuggability of traditional software. Yet serving them at scale remains expensive, largely due to the inflated number of LLM invocations per request. A single workflow execution typically triggers several agent invocations, each making multiple LLM calls; the net effect is that LLM inference can account for nearly 80% of end-to-end request time in workflows (§2.1).
@@
The predominant strategy for reducing these inference costs focuses on intelligently selecting workflow configurations – i.e., end to end assignments of specific LLMs to each agent in the workflow – that minimize cost while preserving accuracy. Existing methods operate at one of two granularities: (1) per-workflow optimizations [cognify, llmselector, murakkab] use historical data to select configurations offline for all inputs, while (2) per-input optimizations [masrouter, routellm, daao] use online routers to select configurations dynamically for each request. Both, however, optimize for static cost metrics (e.g., API dollars or FLOPs) and bind configurations before processing a request, ignoring the increased dynamism that agentic workflows bring to serving systems. Specifically, serving a request is both heterogeneous and lengthy. As it moves through a multi-step workflow graph, the computational cost it brings varies across agents, and the system resources available to it rapidly fluctuate as other requests arrive and execute. Consequently, a configuration selected at the start may quickly become suboptimal during a request’s lifecycle, e.g., an agent assigned a small model to save FLOPs may perform worse than necessary (latency- and throughput-wise) if that model is overloaded and a larger model is idle when the request reaches the agent. Across agentic workloads (§2.2), we find that such static strategies can miss 42.8–233.3% of the potential gains from configuration selection.
To realize the promise of configuration selection, our guiding observation is that the same multi-stage structure that increases the cost of agentic workflows also creates unique opportunities for runtime optimization that existing systems overlook. Workflows exhibit increased fault tolerance – errors made early can often be corrected downstream, e.g., via refinement agents – which broadens the set of viable configurations for each input (configuration flexibility). Further, their stage-by-stage [^1] In this paper, we use the terms stage and agent interchangeably to refer to individual execution steps in an agentic workflow. execution also enables configurations to be adapted as system states evolve, rather than fixed upfront (scheduling flexibility). However, exploiting these flexibilities is challenging. Identifying accurate configurations requires exploring a combinatorial space that grows exponentially with workflow depth and model choices, making per-stage exploration prohibitively expensive. Moreover, stage-wise model decisions must be coordinated across concurrent inputs to account for shared resource contention, further compounding the complexity as more requests and configuration options are introduced.
@@
We present Aragog, a workflow serving system that overcomes these challenges by decomposing configuration selection into two complementary steps. Our key insight is that, for a given input, the accuracy of a configuration is fixed, whereas its runtime performance can change dramatically as system load fluctuates. Aragog leverages this asymmetry by performing the expensive, accuracy-related analysis once per request (i.e., when it arrives) to identify all configurations that preserve accuracy. Then, as the request moves through the workflow, Aragog rapidly selects among only the pre-validated configurations for each upcoming stage, using current runtime signals to stay aligned with system conditions.
@@
To realize this decoupling approach while maximizing serving throughput and avoiding any added latency from configuration decisions, Aragog embeds two optimized components. First, to identify the set of accuracy-preserving configurations, Aragog decomposes the routing problem into many simple binary classification tests and exploits the near-monotonic relationship between configuration FLOPs and accuracy to prune most candidates quickly; this keeps routing both accurate and lightweight enough to overlap with request queuing delays. Second, Aragog’s (stage-wise) runtime scheduler leverages how FIFO priorities and pruned configuration sets naturally structure the assignment space: earlier requests in the queue have only a few viable configurations – since each choice at an earlier stage can invalidate options that use a different model for that stage – and once chosen, they constrain later requests rather than letting the space fan out. This, in turn, creates a compact search space that beam search can traverse efficiently, avoiding both exhaustive exploration and greedy myopia.
@@
We evaluate Aragog by comparing with existing configuration optimizers – both per-workflow [cognify, llmselector, murakkab] and per-input [routellm, masrouter, daao] – augmented with oracle accuracy information about configurations, i.e., representing their best possible performance. Across five popular agentic workflows and multiple model families (Qwen 2.5 [qwen], Llama 3 [llama3], and Phi 4 [phi4]), we find that Aragog improves maximum serving throughput by 50.0–76.3% over per-input optimizations and 78.1–217.0% over per-workflow optimizations. Further, Aragog consistently reduces latency across varying request rates. For instance, at peak load, median and P95 reductions were 32.5–58.8% and 46.8–66.1% over per-input optimizations, and 60.0–78.9% and 63.2–81.5% over per-workflow optimizations. Crucially, Aragog achieves these performance wins while maintaining accuracy within 1% of that when always using the most expensive configuration.
@@ 2 Background and Motivation
We begin by providing an overview of agentic workflows (§2.1), highlighting the limitations of current optimization approaches (§2.2). We then explore the promising opportunities from capitalizing on the unique runtime flexibility of agentic workflows and discuss the practical challenges in effectively realizing these opportunities.
@@ 2.1 Agentic Workflows
Agentic workflows are directed graphs of agent invocations, where each agent—a prompted LLM augmented with tools, memory, and planning capabilities—executes a well-scoped sub-task and passes its output to downstream agents. This structured approach offers similar observability and debuggability expected from conventional software while leveraging LLMs’ generative capabilities—advantages that fully autonomous agents cannot provide [multiagentllmsystemsfail]. As shown in Figure 1, example patterns and application domains include: (1) self-refinement workflows [self_refine] that iteratively improve outputs through generation, critique, and refinement agent; (2) natural language to SQL systems [chess, hexgentext2sql] that decompose query construction into keyword extraction, column selection, SQL generation, and refinement stages; (3) task decomposition workflows [decompose] that split complex problems into parallel sub-tasks handled by different agents before aggregating results; and (4) multi-agent voting systems [self-consistency, voting] that ensemble multiple agents for robust decision-making. In practice, these workflows are explicitly encoded using frameworks like LangChain [langchain] and DSPy [dspy]. Once a workflow graph is defined, developers can specify a configuration that assigns each agent to a chosen LLM, enabling accuracy-cost tradeoffs without modifying workflow logic. These configurations often utilize heterogeneous models [llmselector]–for example, using smaller models for easier agent tasks and larger models for the rest—with each model hosted on different serving engines across different GPUs.
Yet despite these intuitive benefits, agentic workflows face significant practical challenges, largely centered around the high costs they bring. Indeed, each agent invocation incurs the inference cost of powerful yet resource-intensive LLM calls. These expenses scale quickly since workflows generally involve multiple sequential or parallel agent calls. For instance, coding workflows require 3–5 agent calls per query to plan, generate, test, and verify solutions [zhang2023agents, yang2024sweagent], while even simple workflows such as multi-hop question answering for legal or scientific search can involve 5+ agent calls [lewis2020rag, bai2023constitutional, orca]. Across our workloads, LLM calls account for 79.4% of end-to-end latency on average, with tool calls and framework overhead accounting for the remaining 20.6%.
@@ 2.2 Limitations of Existing Optimizations
Existing systems focus on configuration selection to reduce the core LLM inference costs while preserving accuracy, operating at per-workflow or per-input granularity. We categorize and discuss these techniques, highlighting their limitations; complementary optimizations are covered in (§6).
@@
Per-input optimization. In single-LLM serving, existing LLM routing approaches [routellm, mixllm] leverage input heterogeneity by routing it to the most cost-efficient model that meets accuracy requirements. Recent works extend this routing paradigm to agentic workflows [daao, masrouter] by selecting not just singular models, but a workflow configuration for each input. Routers (e.g., transformer-based) are trained using accuracy labels relative to the most expensive configuration, enabling cost reduction while preserving accuracy.
The problem. Though effective, existing optimizations–targeting per-workflow or per-input granularity fail to adapt to the runtime dynamics (e.g., resource availability). The core issue is that they bind configurations before execution and select configurations based on static cost metrics (e.g., FLOPs). As a result, configurations become stale by execution. Worse, the long and multi-stage nature of agentic workflows exacerbates this staleness, as system states can shift dramatically throughout their execution.
@@
Binding too early to configuration selections. Existing approaches all make static configuration decisions prior to any workflow execution, i.e., either once before all inputs, or at the start of each input. The core issue is that system states change rapidly, rendering these a priori selections stale by execution time. Natural LLM dynamics already introduce significant variance: output lengths vary unpredictably, queue depths fluctuate continuously, and dynamic batching reshapes execution efficiency moment by moment [orca]. Agentic workflows amplify this unpredictability by compounding variance across multiple LLM calls. Consequently, any upfront configuration selection before execution can quickly become stale by the time it executes.
@@
Optimizing for static metrics. Current systems minimize predetermined costs like FLOPs or API call prices, but these fail to translate to actual runtime costs, i.e., on metrics such as resource utilization and latency. Instead, what truly governs these aspects is how configurations fit into the current system states and utilize the currently available resources. For instance, a “more expensive” 70B-parameter model on an idle GPU can return results faster than a “cheaper” 14B model with a saturated request queue.
Figure 2: Static approaches (per-workflow, per-input) achieve suboptimal maximal serving capacity. Incorporating runtime cost metrics and runtime configuration adaptation incrementally improves serving capacity. Results are evaluated on HumanEval [humaneval], Bird [bird], and StrategyQA [strategyQA].
To quantify the inefficiency, we follow the experiment setup from (§5.1) and evaluate three workflows–self-refine, text-to-sql, and task decomposition–with different datasets in Figure 1 using three model candidates: Qwen {7,14,32}B [qwen]. For space, we primarily use these workflows here, but note that the trend holds for other workflows and workloads used in our evaluation. In Figure 2, we compare serving capacity (maximum throughput, normalized to the best approach) across configuration selection strategies on three workloads. All approaches are given perfect accuracy knowledge to isolate their fundamental limitations. Per-workflow optimization achieves only 30–45% of maximum capacity. Per-input optimization, which selects the lowest FLOPs accurate configuration for each input, improves serving capacity to 54–63%, demonstrating the benefit of per-input adaptability. Per-input++ is further augmented with a runtime cost oracle that incorporates runtime system metrics into its cost estimation. This improves serving capacity to 70-75%, demonstrating the benefits over using static cost metrics. Per-input++++ further adds a runtime configuration adaptation oracle, adapting configuration based on fresh system states rather than fixing configurations throughout execution. This achieves the highest serving capacity, leaving a 25–30% gap over Per-input++. This significant gap highlights the benefits of continuous and rapid configuration adaptation at runtime over static, upfront configuration selection.
Despite inflating configuration staleness for existing optimizations, the multi-stage structure of agentic workflows also provides unique opportunities to achieve runtime configuration adaptation. Concretely, we identify two forms of flexibility inherent to agentic workflows that enable adaptation to rapid system dynamics.
Configuration Flexibility: Despite the additional complexity and overhead it brings to configuration selection, the exponentially large search space for agentic workflows also presents remarkable error tolerance–when upstream agents produce suboptimal outputs, downstream agents can recover from these intermediate errors (Figure 3). This self-correcting behavior dramatically expands the viable configuration space: for each input, numerous configurations can achieve the target accuracy, not just one “optimal” choice (Figure 4). This large candidate space makes it possible to cater to rapid system dynamics and achieve per-input configuration selection that aligns with dynamic runtime system state, bypassing the limitations of existing routers trained with static cost and upfront fixed configurations.
Scheduling Flexibility: The multi-stage structure of workflows enables dynamic configuration switching during execution. Unlike traditional single LLM systems that have to commit to full configurations upfront, the multi-stage structure naturally allows configuration selection to evolve as the workflow progresses. Specifically, this stage-by-stage execution makes it possible for runtime reconfiguration, which can switch between configurations that share the same prefix–completed stages remain valid while downstream agents can be reassigned based on current system states.
These two forms of flexibility–multiple viable configurations per input and progressive scheduling–highlight the untapped potential in existing systems to adapt configuration to runtime dynamics. However, exploiting this flexibility requires solving fundamental systems challenges in managing the runtime overheads, given the large configuration space and the need for frequent runtime reconfiguration.
@@
Figure 3: Agentic workflows exhibit robustness by recovering from intermediate errors. For each workflow configuration, we track the inputs with intermediate errors, and record the fraction subsequently corrected by the remained workflow. The plot shows the CDF of these correction percentages across all configurations.
Figure 4: Numerous workflow configurations can yield accurate results at per-input granularity.
[…]

### v2

@@ Abstract
Agentic workflows have emerged as a powerful paradigm for solving complex, multi-stage tasks, but serving them at scale is computationally expensive given the many LLM inferences that each request must pass through. Configuration selection, or the cost-aware assignment of workflow agents to specific LLMs, can reduce these costs, but existing approaches bind configuration decisions before request execution, making them ill-suited for the heterogeneous and lengthy execution of workflows. Specifically, system loads can fluctuate rapidly and substantially during a request’s lifetime, causing fixed configurations to quickly become suboptimal. We present Aragog, a system that progressively adapts a request’s configuration throughout its execution to match runtime dynamics. To make this practical despite the massive space of workflow configurations, Aragog decouples the problem into two core elements – a one-time routing step that identifies all accuracy-preserving configurations, and a cheap per-stage scheduler that selects among them using up-to-date system observations – and introduces novel strategies to accelerate each. Across diverse workflows and model families, Aragog increases maximum serving throughput by 42.8–217.0% and reduces median latency by 32.5–86.1% at peak request rates, while maintaining accuracy comparable to the most expensive configurations.
@@ 1 Introduction
Autonomous agents, or Large Language Models (LLMs) augmented with tools, memory, and planning, have shown impressive ability to tackle complex tasks autonomously [yang2024sweagent, snowflake-agentic-workflow]. However, fully autonomous agentic systems often lack the structure and reliability needed in production [multiagentllmsystemsfail]. This has led to the rise of agentic workflows: directed acyclic graphs with fixed structures that orchestrate multiple agents [langchain, dspy, autogen, metagpt], each responsible for a well-defined subtask and passing outputs downstream (Figure 1). These workflows preserve the generative strengths of agents while offering the observability and debuggability of traditional software. Yet serving them at scale remains expensive, largely due to the inflated number of LLM invocations per request. Indeed, a single workflow execution typically triggers several agent invocations, each making multiple LLM calls; the net effect is that LLM inference can account for nearly 80% of end-to-end request time in workflows (§2.1).
@@
The predominant strategy for reducing these inference costs focuses on intelligently selecting workflow configurations – i.e., assignments of specific LLMs to each agent in the workflow – that minimize cost while preserving accuracy. Existing systems make configuration decisions either periodically offline [cognify, llmselector, murakkab] or once per input at request arrival time [masrouter, routellm, daao]. Regardless, they all commit to configurations before executing a request, focusing on optimizing fixed cost metrics such as API dollars or FLOPs. Unfortunately, such a priori decision making is fundamentally misaligned with the heterogeneous and lengthy nature of workflows: as a request slowly moves through a multi-stage graph, the overall system load it encounters can rapidly fluctuate due to the diverse overheads each agent imposes, and the arrival/departure of other requests. Consequently, a configuration selected at the start may quickly become suboptimal during a request’s lifecycle, e.g., an agent assigned a small model to save FLOPs may perform worse than necessary (latency- and throughput-wise) if that model is overloaded and a larger model is idle when the request reaches the agent.
To make matters worse, this early-binding behavior fails to fully exploit the enhanced configuration flexibility that workflows naturally afford through multi-stage processing. Specifically, we observe that workflows exhibit substantial fault tolerance – errors introduced at intermediate stages can often be corrected downstream, e.g., via refinement agents. This, in turn, dramatically expands the set of configurations that preserve end-to-end accuracy, providing a means to cope with (and support) the diverse runtime scenarios that could be encountered during serving. Yet, by binding configurations early, existing approaches forego such opportunities, leaving 25-70% of the potential gains from configuration adaptation on the table for diverse agentic workloads (§2.2).
@@
We present Aragog, an agentic workflow serving system that capitalizes on the aforementioned flexibility by performing stage-wise configuration adaptation. [^1] In this paper, we use the terms stage and agent interchangeably to refer to individual execution steps in an agentic workflow. The primary challenge is one of runtime overheads: per-stage adaptation is prohibitively expensive because it requires frequently exploring an exponentially large configuration space that scales with workflow size and model options. To handle this, our key insight is that, for a given input, the accuracy of a configuration is fixed, whereas its runtime performance can change dramatically as system load fluctuates. Aragog leverages this asymmetry by decomposing configuration selection into two complementary steps. It first performs the expensive, accuracy-related analysis once per request (i.e., when it arrives) to identify all configurations that preserve accuracy. Then, as the request moves through the workflow, Aragog rapidly selects among only the pre-validated configurations for each upcoming stage, using current runtime signals to stay aligned with system conditions.
@@
To realize this decoupling approach while maximizing serving throughput and avoiding any added latency from configuration adaptation, Aragog embeds two optimized components. First, to identify the set of accuracy-preserving configurations, Aragog decomposes the routing problem into many simple binary classification problems and exploits the near-monotonic relationship between configuration FLOPs and accuracy to prune configuration space quickly; this keeps routing both accurate and lightweight enough to overlap with request queuing delays. Second, Aragog’s (stage-wise) runtime scheduler leverages how FIFO priorities and pruned configuration sets naturally structure the assignment space: earlier requests in the queue have only a few viable configurations – since each choice at an earlier stage can invalidate options that use a different model for that stage – and once chosen, they constrain later requests rather than letting the space fan out. This, in turn, creates a compact search space that beam search can traverse efficiently, avoiding both exhaustive exploration and greedy myopia.
@@
We evaluate Aragog by comparing with existing configuration optimizers – both per-workflow [cognify, llmselector, murakkab] and per-input [routellm, masrouter, daao] – augmented with oracle accuracy information about configurations, i.e., representing their best possible performance. Across four popular agentic workflows and diverse model families (Qwen 2.5 [qwen], Llama 3 [llama3], and Phi 4 [phi4]), we find that Aragog improves maximum serving throughput by 42.8–76.3% over per-input optimizations and 78.1–217.0% over per-workflow optimizations. Further, Aragog consistently reduces latency across varying request rates. For instance, at peak load, median and P95 reductions were 32.5–71.1% and 46.2–76.2% over per-input optimizations, and 60.0–86.1% and 63.2–89.0% over per-workflow optimizations. Crucially, Aragog achieves these performance wins while maintaining accuracy within 2% of that when always using the most expensive configuration. We will open source Aragog post publication.
@@ 2 Background and Motivation
We begin by providing an overview of agentic workflows (§2.1), highlighting the limitations of current optimization approaches and opportunities they overlook (§2.2). We then discuss the practical challenges in realizing them (§2.3).
@@ 2.1 Agentic Workflows
Agentic workflows are directed graphs of agent invocations, where each agent—a prompted LLM augmented with tools, memory, and planning capabilities—executes a well-scoped sub-task and passes its output to downstream agents. As shown in Figure 1, example patterns and application domains include: (1) self-refine workflows [self_refine, llmselector] that iteratively improve outputs through generation, critique, and refinement agents; (2) natural language to SQL systems [chess, hexgentext2sql] that decompose query construction into keyword extraction, column selection, SQL generation, and refinement; (3) task decomposition workflows [decompose, autogen] that split complex problems into parallel sub-tasks handled by different agents before aggregating their results; and (4) multi-agent voting systems [self-consistency, voting] that ensemble multiple agents for robust decision-making. In practice, these workflows are explicitly encoded using frameworks like LangChain [langchain] and DSPy [dspy]. Once a workflow graph is defined, developers can specify a configuration that assigns each agent to a chosen LLM, enabling accuracy-cost tradeoffs without modifying workflow logic. These configurations often utilize heterogeneous models – i.e., smaller models for easier agent tasks, specialized models for domain-specific tasks – that are each hosted on different serving engines and GPUs [llmselector].
Yet despite these intuitive benefits, agentic workflows face significant practical challenges, largely centered around the high costs they bring. Indeed, each agent invocation incurs the inference cost of powerful yet resource-intensive LLM calls. These expenses scale quickly since workflows generally involve multiple sequential or parallel agent calls. For instance, coding workflows require 3–5 agent calls per query to plan, generate, test, and verify solutions [zhang2023agents, yang2024sweagent], while even simple workflows such as multi-hop question answering for legal or scientific search can involve 4+ agent calls [lewis2020rag, bai2023constitutional, orca]. Across our workloads, a request invokes 5.2 LLM calls on average. These LLM calls account for 79.4% of end-to-end latency on average, with tool calls and framework (e.g., DSPy) overhead accounting for the remaining 20.6%.
@@ 2.2 Limitations of Existing Optimizations
Existing systems primarily focus on configuration selection to reduce these LLM inference costs while preserving workflow accuracy. We categorize and discuss these techniques below, highlighting their limitations; complementary optimizations are covered in §6.
@@
Per-input optimization. In single-LLM serving, existing LLM routing approaches [routellm, mixllm] leverage input heterogeneity by training routers to forward each input to the most cost-efficient model that can accurately respond to that input. Recent works extend this routing paradigm to agentic workflows [daao, masrouter] by selecting not just singular models, but a workflow configuration for each input. Routers are often transformer-based and are trained using accuracy labels relative to the most expensive configuration, enabling cost reduction while preserving accuracy.
@@
The problem. Though effective, existing optimizations fail to adapt to the intrinsic runtime dynamics of workflows, falling short in two key ways. First, they bind configurations before workflow execution, i.e., either once before all inputs, or at the start of each input. The core issue is that as a request is executing, system states can change rapidly, rendering these a priori selections suboptimal by execution time. Natural LLM dynamics already introduce significant variance: output lengths vary unpredictably, queue depths fluctuate continuously, and dynamic batching reshapes execution efficiency moment by moment [orca]. The lengthy, multi-stage nature of agentic workflows only exacerbates this; the load for each concurrent request can change dramatically as it shifts between workflow agents with different overheads (and potentially different models), and the set of concurrent requests can fluctuate many times as a request slowly passes through the entire workflow graph. Consequently, upfront configuration selections can increasingly become suboptimal as requests progress through multi-stage execution.
@@
Second, these same dynamics make the approach of optimizing for static cost metrics problematic. Indeed, current systems all minimize predetermined costs like FLOPs or API call prices, but these commonly fail to translate to actual runtime costs (and thus, performance goals), i.e., for metrics such as throughput, resource utilization, and latency. Instead, what truly governs these aspects is how configurations fit into the current system states and utilize the currently available resources. For instance, a “more expensive” 70B-parameter model on an idle GPU can return results faster than a “cheaper” 14B model with a saturated request queue.
@@
Figure 2: Agentic workflows exhibit robustness by recovering from intermediate errors. For each configuration, we track the inputs with intermediate errors, and record the fraction subsequently corrected by the remained workflow. The plot shows the CDF of correction percentages across all configurations.
Figure 3: Numerous workflow configurations can match the accuracy of the most expensive configuration at for each input.
These drawbacks are even more problematic when contextualized relative to the unique opportunities that workflows bring. Most notably, despite the increased LLM inference overheads, the multi-stage execution of workflows provides remarkable configuration flexibility that exceeds what is typically possible with single-model pipelines, e.g., with classic model routing. Concretely, given multiple model options per stage, agentic workflows elicit an exponentially large configuration space that presents remarkable error tolerance – when upstream agents produce suboptimal outputs, downstream agents can often recover from those intermediate errors (Figure 2). This self-correcting behavior dramatically expands the viable configuration space (Figure 3), highlighting an untapped potential: the enhanced configuration flexibility provides support for frequently adapting to dynamic runtime scenarios via stage-wise configuration adaptation using the many options. By making static configuration decisions prior to workflow execution, all existing approaches fail to fully take advantage of this enhanced flexibility.
Figure 4: Static approaches (per-workflow, per-input) achieve suboptimal serving capacity. Incorporating runtime cost metrics and runtime configuration adaptation incrementally improves serving capacity. Results are evaluated on HumanEval [humaneval], Bird [bird], and StrategyQA [strategyQA].
To quantify this potential (and the drawbacks of existing approaches), we follow the setup from §5.1 and evaluate three representative workflows—self refine, text-to-sql, and task decomposition—with their corresponding datasets (HumanEval [humaneval], Bird [bird], StrategyQA [strategyQA]) and Qwen {7,14,32}B as model options. We compare the serving capacity (max achieved throughput, normalized to the best approach) of Per-workflow and Per-input configuration strategies (as described above), as well as two more schemes:
  • $\bullet$
    Per-input++, which improves upon Per-input by factoring in runtime observations (in particular, the predicted latency based on current load for each model and serving engine) rather than solely FLOPs when selecting a configuration prior to each request, and
  • $\bullet$
    Per-input++++, which improves Per-input++ by continuously re-evaluating and updating configuration selections at each workflow stage (rather than solely before each request) to adapt to changing system dynamics.
To isolate impact of selection strategy, all approaches are given perfect knowledge of accuracy per configuration.
As shown in Figure 4, despite the benefits that Per-input strategies bring via input-level configuration adaptation (i.e., 9–33% improvement in serving capacity over Per-workflow schemes), their early binding and static cost metrics leave substantial performance gains on the table. In particular, our results highlight that incorporating dynamic cost metrics (Per-input++) and stage-wise configuration adaptation (Per-input++++) can further improve serving capacity by 12–16% and 25–30% when they are incrementally incorporated – a 37–46% increase in maximum achieved throughput over even the best possible version of existing Per-input schemes.
2.3 Challenges
Despite the benefits, realizing stage-wise configuration adaptation requires solving fundamental systems challenges in managing runtime overheads, given the large configuration space to explore and the need for frequent reconfiguration.
C1: Prohibitive and frequent routing overheads. Agentic workflows create configuration spaces that grow exponentially with workflow complexity: $M^{N}$ configurations for $N$ agents and $M$ models. Routing must evaluate this massive space to select configurations for each input. Yet, routing overhead scales poorly with the configuration space: state-of-the-art routers already build on large transformers for accurate single-LLM selection [routellm, mixllm], and workflow routing demands substantially more complexity to evaluate exponentially more candidates. Moreover, adapting to rapid system dynamics requires frequent rerouting for every input. The untenable and frequent routing cost for every input can offset the benefit of flexible configuration adaptation.